Solve the system;
x+y=20
x-y=8

Answer by checkley71(8403)

X+Y=20
X-Y=8 NOW ADD THESE 2 EQUATIONS.
-------------------
2X=28
X=28/2
X=14 ANSWER.
14+Y=20
Y=20-14
Y=6 ANSWER.
PROOF
14-6=8
8=8